• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 04:32:56 on July29,2003

    Entry number 107863

    Followups:

    keyword=DAQ problems/solutions

    Since I could not diagnose the DAQ problems from home I came in.

    On the R-arm I rebooted everything from scratch and it works fine again.
    The only thing I did differently is that I made sure the ET system was killed off.
    Oddly I could not kill it from the a-onl account so I used "root" to "kill -9".

    On the L-arm the problem is two-fold: 1) The normalization scalers (both ungated
    by helicity and gated by helicity) read back zero. 2) Deadtime is very high, e.g.
    ~250 Hz observed rate for ~500 Hz pulser (--> 50% DT). This is probably related
    to the scaler problem (problem #1) if there is a timeout for readout. I tried
    rebooting the trigger supervisor crate (TS1) several times. The solution should
    be to power cycle the TS1 crate. The experiment doesn't want to take access on L-arm
    now, so L-arm DAQ is presently running with no scalers and high DT.